OpenStudy (anonymous):

How do you simplify 3/7?

OpenStudy (anonymous):

3 divided by 7

OpenStudy (jdoe0001):

well... I don't think you can, is as simplified as it can be

OpenStudy (jdoe0001):

beating up a dead cat doesn't make it deader, just flatter

OpenStudy (anonymous):

hahahahahaha!!! so true dude! im going to fan you! , can you help me with m y question?

OpenStudy (jdoe0001):

dunno.... what is it?

OpenStudy (anonymous):

Nick wants to buy a pair of shoes. The original cost of the shoes is $56.75, and the markup is 12 percent. How much will he have to pay for the shoes? markup value = original cost × percentage markup Select one of the options below as your answer: A. $6.81 B. $49.64 C. $63.56

OpenStudy (jdoe0001):

well, he'd have to pay for the shoes, 56.75 + tax, that is 56.75 + (12% of 56.75) \(\bf \textit{n% of m} = \cfrac{n}{100}\times m\\ \quad \\ \textit{so the full price will be} \qquad 56.75 + \left(\cfrac{12}{100}\times 56.75\right)\)